OpenStudy (anonymous):

(g + 4) – 3g = 1 + g (1 point)A.1 B.4 C.no solution D.identity

OpenStudy (anonymous):

c

OpenStudy (anonymous):

c

OpenStudy (anonymous):

(g + 4) – 3g = 1 + g (1 point) 1 4 no solution identity 2. 5y + 2 = (10y + 4) (1 point) 5 10 no solution identity 3. 0.5b + 4 = 2(b + 2) (1 point) 0 0.5 no solution identity 4. 8 – (3 + b) = b – 9 (1 point) 7 8 no solution identity

OpenStudy (anonymous):

welcome